OpenStudy (anonymous):

f(x) = 1/(1-x^2) through x^8 term is = X^8+x^6+x^4+x^2+1, use partial expansion to in power of x to come up with the value of derivative f'(0),f''(0),f'''(0)...f''''''''(0)

OpenStudy (dumbcow):

what do you mean by "through x^8 term" ? f(x) = 1/(1-x^2) correct ?

OpenStudy (anonymous):

in 8 times power

OpenStudy (anonymous):

grind it til you find it \[f'(x)=8x^7+6x^5+4x^3+2x\implies f'(0)=0\]

OpenStudy (anonymous):

\[f''(0)=2\] etc

OpenStudy (anonymous):

so f'''(0)=0 and f''''(0)=4 ?

OpenStudy (anonymous):

\(f'''(0)=0\) but i don't think \(f^{(4)}(0)=4\)

OpenStudy (anonymous):

why? i though in power of 4 times, the number is 4?

OpenStudy (anonymous):

i think it is 24 but it is late and i am doing it in my head try and see what you get

OpenStudy (anonymous):

how did you get the 24? how did you calculate

OpenStudy (anonymous):

it is from the term that started out at \(x^4\) when you take successive derivatives you get \(4x^3, 12x^2, 24x, 24\) aka 4!

OpenStudy (anonymous):

so \(f^{(4)}(0)=4!\) which i think is where this is headed and of course \(f^{(5)}(0)=0\)

OpenStudy (anonymous):

want to guess at \(f^{(6)}(0)\) ?

OpenStudy (anonymous):

is 6!?

OpenStudy (anonymous):

you found the sound

OpenStudy (anonymous):

it is only apply for this formula or it's universal?

OpenStudy (anonymous):

of course you can check by differentiating, but i am sure that is what you will get

OpenStudy (anonymous):

well you have coefficients of 1 for each term, so that makes it easier also this is the derivative at zero, so we have to be careful when generalizing if the expansion was different, you would get different answers, but there would certainly be an \(n!\) involved, as you can see

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you are being set up to learn about taylor polynomials, where the coefficients look like \[\frac{f^{(n)}(0)}{n!}\] yw

OpenStudy (anonymous):

if a is not 0, just an "a", where we should put (x-a), after n! or other place?